OpenStudy (anonymous):

what is the sum of the roots of the polynomial shown below? f(x)=x^3-8x^2-23x+30

OpenStudy (anonymous):

sum of roots=-b/a comparing it with standard form i.e. ax^3+bx^2..... b=-8 and a=1 hence sum of roots =-8/1 =-8

OpenStudy (anonymous):

Oh! Except that -b would make b=+8, but i gotcha! Thanks a lot c:

OpenStudy (anonymous):

good